SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (12th Dec, 2006)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Fang Fang, Rahmat Rahmat, Su Dong.

Operations

  • The spares situation is under control, but should check there's a full set in the IR2 spares cabinet - useless boards (TSFY etc) can be moved to the lab.
  • The priority for fixing are the BLTs - both spares need to be checked if they have the new firmware or not. Su Dong will help Debbie with this. The 'bad' DCCs need to be diagnosed in the lab, and then sent to France for repairs. The bad FCPM can be given to the dataflow people and the TSFi and ZPDis should be fixed. Old TSF and PTD boards are no longer needed so they should be `stored' somewhere. Once fixed, spares should rotated into the system.
  • Cosmics run for DCH expected 18th/19th Dec. During this, can check the IFS timing is correct. We need to make sure we import the DCH calibration into L3 once it's done and check the effect of the higher coltage - may need to re-calibrate.

EMT

  • No word from Dan about mumu plots yet.
  • Histograms from pre- and post-fixed L1EmtSim have been compared and look identical, so Mark will go ahead with comitting the new package.

L3

  • Arik sent 2 xtc and 2 moose collections (with the old and new L3 configurations) to the tracking group who will validate the changes.

Data Quality and Monitoring

  • A couple of possibly dodgy runs will be looked at by the DQM group on their Thursday meeting.
  • Rahmat had some problems with root hanging on one run when making the Run 5a stripcharts. He will look into this.

Config

  • See here for meeting notes.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.